# AdaRC: Mitigating Graph Structure Shifts during Test-Time

### v1

@@ Abstract
Powerful as they are, graph neural networks (GNNs) are known to be vulnerable to distribution shifts. Recently, test-time adaptation (TTA) has attracted attention due to its ability to adapt a pre-trained model to a target domain, without re-accessing the source domain. However, existing TTA algorithms are primarily designed for attribute shifts in vision tasks, where samples are independent. These methods perform poorly on graph data that experience structure shifts, where node connectivity differs between source and target graphs. We attribute this performance gap to the distinct impact of node attribute shifts versus graph structure shifts: the latter significantly degrades the quality of node representations and blurs the boundaries between different node categories. To address structure shifts in graphs, we propose AdaRC, an innovative framework designed for effective and efficient adaptation to structure shifts by adjusting the hop-aggregation parameters in GNNs. To enhance the representation quality, we design a prediction-informed clustering loss to encourage the formation of distinct clusters for different node categories. Additionally, AdaRC seamlessly integrates with existing TTA algorithms, allowing it to handle attribute shifts effectively while improving overall performance under combined structure and attribute shifts. We validate the effectiveness of AdaRC on both synthetic and real-world datasets, demonstrating its robustness across various combinations of structure and attribute shifts.
@@ 1 Introduction
Graph neural networks (GNNs) have shown great success in various graph applications such as social networks [31], scientific literature networks [12], and financial fraud detection [28]. Their superior performance heavily relies on the assumption that training and testing graph data are identically distributed [18]. However, real-world graphs usually involve distribution shifts in both node attributes and graph structures [23, 38, 39]. For example, given two social networks (e.g., LinkedIn for professional networking and Pinterest for casual content sharing), the user profiles are likely to vary due to the different functionalities of two graphs, resulting in attribute shifts. Besides, as LinkedIn users tend to connect with professional colleges, while users on Pinterest often connect with family and friends, the connectivity patterns vary across different networks, introducing structure shifts. The co-existence of these complex shifts significantly undermines GNN model performance [18].
Various approaches have been proposed to address the distribution shifts between the source and target domains, e.g., domain adaptation [35] and domain generalization [34]. But most of these approaches require access to either target labels [38, 39] or the source domain during adaptation [23, 42], which is often impractical in real-world applications. For example, when a model is deployed for fraud detection, the original transaction data used for training may no longer be accessible. Test-time adaptation (TTA) has emerged as a promising solution, allowing models to adapt to an unlabeled target domain without re-accessing the source domain [21]. These algorithms demonstrate robustness against various image corruptions and style shifts in vision tasks [33, 13, 47]. However, when applied to graph data, existing TTA algorithms face significant challenges, especially under structure shifts. As shown in Figure 2, both attribute and structure shifts (e.g., homophily and degree shifts) lead to performance drops on target graphs, but current TTA methods provide only marginal accuracy improvements under structure shifts compared to attribute shifts.
Figure 1: Generic TTA algorithms (T3A, Tent, AdaNPC) are significantly less effective under structure shifts (right) than attribute shifts (left). On the contrary, our proposed AdaRC could significantly improve the performance of generic TTA (gray shaded area). The dataset used is CSBM.
@@
In this paper, we seek to understand why generic TTA algorithms perform poorly under structure shifts. Through both theoretical analysis and empirical evaluation, we reveal that while both attribute and structure shifts affect model accuracy, they impact GNNs in different ways. Attribute shifts mainly affect the decision boundary and can often be addressed by adapting the downstream classifier. In contrast, structure shifts degrade the upstream featurizer, causing node representations to mix and become less distinguishable, which significantly hampers performance. Figure 2 illustrates this distinction. Since most generic TTA algorithms rely on high-quality representations [13, 33, 47], they struggle to improve GNN performance under structure shifts.
To address these limitations, we propose that the key to mitigating structure shifts lies in restoring the quality of node representations, making the representations of different classes distinct again. Guided by theoretical insights, we propose adjusting the hop-aggregation parameters which control how GNNs integrate node features with neighbor information across different hops. Many GNN designs include such hop-aggregation parameters, e.g., GPRGNN [7], APPNP [17], JKNet [43], and GCNII [6]. Building on this, we introduce our AdaRC framework. It restores representation quality impacted by structure shifts by adapting hop-aggregation parameters through minimizing prediction-informed clustering (PIC) loss, promoting discriminative node representations without falling into trivial solutions as with traditional entropy loss. Additionally, our framework can be seamlessly integated with existing TTA algorithms to harness their capability to handle attribute shifts. We empirically evaluate AdaRC with a wide range of datasets and TTA algorithms. Extensive experiments on both synthetic and real-world datasets show that AdaRC can handle a variety of structure shifts, including homophily shifts and degree shifts. Moreover, it is compatible to a wide range of TTA algorithms and is able to enhance their performance under various combinations of attribute shifts and structure shifts. We summarize our contributions as follows:
@@
  • A novel framework AdaRC is proposed to restore the quality of node representations and boost existing TTA algorithms by adjusting the hop-aggregation parameters.
  • Empirical evaluation on both synthetic and real-world scenarios demonstrates the effectiveness of AdaRC under various distribution shifts. When applied alone, AdaRC enhances the source model performance by up to 31.95%. When integrated with existing TTA methods, AdaRC further boosts their performance by up to 40.61%.
@@ 2 Related works
Test-time adaptation (TTA) aims to adapt a pre-trained model from the source domain to an unlabeled target domain without re-accessing the source domain during adaptation [21]. For i.i.d. data like images, several recent works propose to perform image TTA by entropy minimization [33, 46], pseudo-labeling [13, 47], consistency regularization [3], etc. However, graph TTA is more challenging due to the co-existence of attribute shifts and structure shifts. To address this issue, GTrans [14] proposes to refine the target graph at test time by minimizing a surrogate loss. SOGA [27] maximizes the mutual information between model inputs and outputs, and encourages consistency between neighboring or structurally similar nodes, but it is only applicable to homophilic graphs. Focusing on degree shift, GraphPatcher [15] learns to generate virtual nodes to improve the prediction on low-degree nodes. In addition, GAPGC [5] and GT3 [37] follow a self-supervised learning (SSL) scheme to fine-tune the pre-trained model for graph classification.
Graph domain adaptation (GDA) aims to transfer knowledge from a labeled source graph to an unlabeled target graph with access to both graphs. Most of the GDA algorithms focus on learning invariant representations over the source and target graphs by adversarial learning [48, 41, 42] or minimizing the distance between source and target graphs [52, 39]. More recent work [23] addresses the co-existence of structure and node attribute shifts by reweighing the edge weights of the target graphs. However, GDA methods require simultaneous access to both the source and target graphs, and thus cannot be extended to TTA scenarios.
@@ 3.1 Preliminaries
In this paper, we focus on graph test-time adaptation (GTTA) for node classification. A labeled source graph is denoted as ${\mathcal{S}}=({\bm{X}}_{S},{\bm{A}}_{S})$ with node attribute matrix ${\bm{X}}_{S}\in\mathbb{R}^{N\times D}$ and adjacency matrix ${\bm{A}}_{S}\in\{0,1\}^{N\times N}$. The corresponding node label matrix is denoted as ${\bm{Y}}_{S}\in\{0,1\}^{N\times C}$. For a node $v_{i}$, we denote its neighbors as ${\mathbb{N}}(v_{i})$ and node degree as $d_{i}$. A GNN model $g_{S}\circ f_{S}(\cdot)$ is pre-trained on the source graph, where $f_{S}$ is the featurizer extracting node-level representations, and $g_{S}$ is the classifier, which is usually a linear layer. The goal of GTTA is to adapt the pre-trained GNN model to enhance node classification accuracy on an unlabeled target graph ${\mathcal{T}}=({\bm{X}}_{T},{\bm{A}}_{T})$ with a different distribution, while the source graph ${\mathcal{S}}$ are not accessible during adaptation. [^1] This setting is also referred to as source-free unsupervised graph domain adaptation [27]. Here, we primarily follow the terminology used in [14]. It is important to note that, unlike the online setting often adopted in image TTA, graph TTA allows simultaneous access to the entire unlabeled target graph ${\mathcal{T}}$ [21].
Compared with TTA on regular data like images, GTTA is more challenging due to the co-existence of attribute shifts and structure shifts [18, 39], which are formally defined as follows [23].
@@
Homophily refers to the phenomenon that a node tends to connect with nodes with the same labels. Formally, the node homophily of a graph $\mathcal{G}$ is defined as [30]:
$h(\mathcal{G})=\frac{1}{N}\!\sum_{i}h_{i},\quad\text{where }h_{i}=\frac{|\{u\!\in\!{\mathbb{N}}(v_{i}):y_{u}\!=\!y_{v}\}|}{d_{i}},$ (1)
@@ 3.2 Impacts of distribution shifts
CSBM [8] is a random graph generator widely used in the analysis of GNNs [25, 26, 44]. Specifically, we consider a CSBM with two classes +=${\mathbb{C}}_{+}=\{v_{i}:y_{i}=+1\}$ and =${\mathbb{C}}_{-}=\{v_{i}:y_{i}=-1\}$, each having $\frac{N}{2}$ nodes. The attributes for each node $v_{i}$ are independently sampled from a Gaussian distribution ${\bm{x}}_{i}\sim{\mathcal{N}}({\bm{\mu}}_{i},{\bm{I}})$, where ${\bm{\mu}}_{i}={\bm{\mu}}_{+}$ for $v_{i}\in{\mathbb{C}}_{+}$ and ${\bm{\mu}}_{i}={\bm{\mu}}_{-}$ for $v_{i}\in{\mathbb{C}}_{-}$. Each pair of nodes are connected with probability $p$ if they are from the same class, otherwise $q$. As a result, the average degree is $d=\frac{N(p+q)}{2}$ and node homophily is $h=\frac{p}{p+q}$. We denote the graph as $\text{CSBM}({\bm{\mu}}_{+},{\bm{\mu}}_{-},d,h)$, where ${\bm{\mu}}_{+},{\bm{\mu}}_{-}$ encode the node attributes and $d,h$ encode the graph structure.
@@
We consider a single-layer GCN, whose featurizer is denoted as ${\bm{Z}}\!=\!f({\bm{X}},\!{\bm{A}})\!=\!{\bm{X}}+\gamma\cdot{\bm{D}}^{-1}{\bm{A}}{\bm{X}}=({\bm{I}}+\gamma\cdot{\bm{D}}^{-1}{\bm{A}}){\bm{X}}$, where ${\bm{D}}$ is the degree matrix. Equivalently, for each node $v_{i}$, its node representation is ${\bm{z}}_{i}={\bm{x}}_{i}+\gamma\cdot\frac{1}{d_{i}}\sum_{v_{j}\in{\mathcal{N}}(v_{i})}{\bm{x}}_{j}$. The parameter $\gamma$ controls the mixture between the node’s own representation and its one-hop neighbors’ average representation. We consider $\gamma$ as a fixed parameter for now, and adapt it later in Subsection 3.3. We consider a linear classifier as $g({\bm{Z}})={\bm{Z}}{\bm{w}}+{\bm{1}}b$, which predicts a node $v_{i}$ as positive if ${\bm{z}}_{i}^{\top}{\bm{w}}+b\geq 0$ and vise versa.
@@
Proposition 3.1.
For graphs generated by $\text{CSBM}({\bm{\mu}}_{+},{\bm{\mu}}_{-},d,h)$, the node representation ${\bm{z}}_{i}$ of node $v_{i}\in{\mathbb{C}}_{+}$ generated by a single-layer GCN follows a Gaussian distribution of
@@
Corollary 3.2.
When ${\bm{\mu}}_{+}={\bm{\mu}},{\bm{\mu}}_{-}=-{\bm{\mu}}$, and all nodes have the same homophily $h=\frac{p}{p+q}$ and degree $d=\frac{N(p+q)}{2}$, the classifier maximizes the expected accuracy when ${\bm{w}}=\operatorname{sign}(1+\gamma(2h-1))\cdot\frac{{\bm{\mu}}}{\|{\bm{\mu}}\|_{2}}$ and $b=0$. It gives a linear decision boundary of $\{{\bm{z}}:{\bm{z}}^{\top}{\bm{w}}=0\}$ and the expected accuracy
@@
To analyze the distinct impact patterns of attribute shifts and structure shifts, we decompose the accuracy gap of GNNs between the source graph and the target graph into two parts as follows,
@@
where $\text{Acc}_{S},\text{Acc}_{T}$ denote the accuracies on the source and target graphs, respectively. $\sup_{g_{T}}\text{Acc}_{T}(g_{T}\circ f_{S})$ is the highest accuracy that a GNN can achieve on the target graph when the featurizer $f_{S}$ is frozen and the classifier $g_{T}$ is allowed to adapt. Using this accuracy as a pivot, the accuracy gap is decomposed into representation degradation and classifier bias. A visualized illustration is shown in Figure 7.
@@
Proposition 3.3 (Impacts of attribute shifts).
When training a single-layer GCN on a source graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d,h)$, while testing it on a target graph of $\text{CSBM}({\bm{\mu}}+\Delta{\bm{\mu}},-{\bm{\mu}}+\Delta{\bm{\mu}},d,h)$ with $\|\Delta{\bm{\mu}}\|_{2}<|\frac{1+\gamma(2h-1)}{1+\gamma}|\cdot\|{\bm{\mu}}\|_{2}$, we have
@@
Proposition 3.4 (Impacts of structure shifts).
When training a single-layer GCN on a source graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d_{S},h_{S})$, while testing it on a target graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d_{T},h_{T})$, where $1\leq d_{T}=d_{S}-\Delta d<d_{S}$ and $\frac{1}{2}<h_{T}=h_{S}-\Delta h<h_{S}$, if $\gamma>0$, we have
@@
Propositions 3.3 and 3.4 imply that attribute shifts and structure shifts impact the accuracy of GNN differently. Specifically, attribute shifts impact the decision boundary of the classifier, while structure shifts significantly degrade the node representation quality. These propositions also match with our empirical findings in Figure 2 and Figure 9 (in Appendix C.1). Since generic TTA methods [33, 13, 47] usually rely on the representation quality and refine the decision boundary, their effectiveness is limited under structure shifts.

### v2

# Matcha: Mitigating Graph Structure Shifts with Test-Time Adaptation

@@ Abstract
Powerful as they are, graph neural networks (GNNs) are known to be vulnerable to distribution shifts. Recently, test-time adaptation (TTA) has attracted attention due to its ability to adapt a pre-trained model to a target domain, without re-accessing the source domain. However, existing TTA algorithms are primarily designed for attribute shifts in vision tasks, where samples are independent. These methods perform poorly on graph data that experience structure shifts, where node connectivity differs between source and target graphs. We attribute this performance gap to the distinct impact of node attribute shifts versus graph structure shifts: the latter significantly degrades the quality of node representations and blurs the boundaries between different node categories. To address structure shifts in graphs, we propose Matcha, an innovative framework designed for effective and efficient adaptation to structure shifts by adjusting the hop-aggregation parameters in GNNs. To enhance the representation quality, we design a prediction-informed clustering loss to encourage the formation of distinct clusters for different node categories. Additionally, Matcha seamlessly integrates with existing TTA algorithms, allowing it to handle attribute shifts effectively while improving overall performance under combined structure and attribute shifts. We validate the effectiveness of Matcha on both synthetic and real-world datasets, demonstrating its robustness across various combinations of structure and attribute shifts. Our code is available at https://github.com/baowenxuan/Matcha.
@@ 1 Introduction
Graph neural networks (GNNs) have shown great success in various graph applications such as social networks (Rozemberczki et al., 2021), scientific literature networks (Hu et al., 2020), and financial fraud detection (Pareja et al., 2020). Their success heavily relies on the assumption that training and testing graphs are identically distributed (Li et al., 2022a). However, real-world graphs usually involve distribution shifts in both node attributes and graph structures (Liu et al., 2023; Wu et al., 2023a; b). For example, given two social networks (e.g., LinkedIn for professional networking and Instagram for casual content sharing), the user profiles are likely to vary due to the different functionalities of two graphs, resulting in attribute shifts. Besides, as LinkedIn users tend to connect with professional colleges, while users on Instagram often connect with family and friends, the connectivity patterns vary across different networks, introducing structure shifts. The co-existence of these complex shifts significantly undermines GNN model performance (Li et al., 2022a).
Various approaches have been proposed to tackle distribution shifts between the source and target domains, e.g., domain adaptation (Wang & Deng, 2018) and domain generalization (Wang et al., 2023b). But most of these approaches require access to either target labels (Wu et al., 2023a; b) or the source domain during adaptation (Liu et al., 2023; Xiao et al., 2023), which is often impractical in real-world applications. For example, when a model is deployed for fraud detection, the original transaction data used for training may no longer be accessible. Test-time adaptation (TTA) has emerged as a promising solution, allowing models to adapt to an unlabeled target domain without re-accessing the source domain (Liang et al., 2023). These algorithms demonstrate robustness against various image corruptions and style shifts in vision tasks (Wang et al., 2021; Iwasawa & Matsuo, 2021; Zhang et al., 2023). However, applying TTA to graph data presents significant challenges, especially under structure shifts. As shown in Figure 2, both attribute and structure shifts (e.g., homophily and degree shifts) lead to performance drops on target graphs, but current TTA methods provide only marginal accuracy improvements under structure shifts compared to attribute shifts.
Figure 1: Generic TTA algorithms (T3A, Tent, AdaNPC) are significantly less effective under structure shifts (right) than attribute shifts (left). On the contrary, our proposed Matcha could significantly improve the performance of generic TTA (gray shaded area). The dataset used is CSBM.
@@
In this paper, we seek to understand why generic TTA algorithms perform poorly under structure shifts. Through both theoretical analysis and empirical evaluation, we reveal that while both attribute and structure shifts affect model accuracy, they impact GNNs in different ways. Attribute shifts mainly affect the decision boundary and can often be addressed by adapting the downstream classifier. In contrast, structure shifts degrade the upstream featurizer, causing node representations to mix and become less distinguishable, which significantly hampers performance. Figure 2 illustrates this distinction. Since most generic TTA algorithms rely on high-quality representations (Iwasawa & Matsuo, 2021; Wang et al., 2021; Zhang et al., 2023), they struggle to improve GNN performance under structure shifts.
To address these limitations, we propose that the key to mitigating structure shifts lies in restoring the quality of node representations, making the representations of different classes distinct again. Guided by theoretical insights, we propose adjusting the hop-aggregation parameters which control how GNNs integrate node features with neighbor information across different hops. Many GNN designs include such hop-aggregation parameters, e.g., GPRGNN (Chien et al., 2021), APPNP (Klicpera et al., 2019), JKNet (Xu et al., 2018), and GCNII (Chen et al., 2020). Building on this, we introduce Matcha, a framework to Mitigate grAph sTruCture sHifts with test-time Adaptation. It restores representation quality by adapting hop-aggregation parameters via minimizing prediction-informed clustering (PIC) loss, promoting discriminative node representations without falling into trivial solutions as with traditional entropy loss. Additionally, our framework can be seamlessly integated with existing TTA algorithms to harness their capability to handle attribute shifts. We empirically evaluate Matcha with a wide range of datasets and TTA algorithms. Extensive experiments on both synthetic and real-world datasets show that Matcha can handle a variety of structure shifts, including homophily shifts and degree shifts. Moreover, it is compatible to a wide range of TTA algorithms and is able to enhance their performance under various combinations of attribute shifts and structure shifts. We summarize our contributions as follows:
@@
  • A novel framework Matcha is proposed to restore the quality of node representations and boost existing TTA algorithms by adjusting the hop-aggregation parameters.
  • Empirical evaluation on both synthetic and real-world scenarios demonstrates the effectiveness of Matcha under various distribution shifts. When applied alone, Matcha enhances the source model performance by up to 31.95%. When integrated with existing TTA methods, Matcha further boosts their performance by up to 40.61%.
@@ 2 Related works
Test-time adaptation (TTA) aims to adapt a pre-trained model from the source domain to an unlabeled target domain without re-accessing the source domain during adaptation (Liang et al., 2023). For i.i.d. data like images, several recent works propose to perform image TTA by entropy minimization (Wang et al., 2021; Zhang et al., 2022), pseudo-labeling (Iwasawa & Matsuo, 2021; Zhang et al., 2023), consistency regularization (Boudiaf et al., 2022), etc. However, graph TTA is more challenging due to the co-existence of attribute shifts and structure shifts. To address this issue, GTrans (Jin et al., 2023) proposes to refine the target graph at test time by minimizing a surrogate loss. SOGA (Mao et al., 2024) maximizes the mutual information between model inputs and outputs, and encourages consistency between neighboring or structurally similar nodes, but it is only applicable to homophilic graphs. Focusing on degree shift, GraphPatcher (Ju et al., 2023) learns to generate virtual nodes to improve the prediction on low-degree nodes. In addition, GAPGC (Chen et al., 2022) and GT3 (Wang et al., 2022) follow a self-supervised learning (SSL) scheme to fine-tune the pre-trained model for graph classification.
Graph domain adaptation (GDA) aims to transfer knowledge from a labeled source graph to an unlabeled target graph with access to both graphs. Most of the GDA algorithms focus on learning invariant representations over the source and target graphs by adversarial learning (Zhang et al., 2019; Wu et al., 2020; Xiao et al., 2023) or minimizing the distance between source and target graphs (Zhu et al., 2021b; Wu et al., 2023b). More recent works (Liu et al., 2023; 2024a) address the co-existence of structure and node attribute shifts by reweighing the edge weights of the source graphs. However, GDA methods require simultaneous access to both the source and target graphs, and thus cannot be extended to TTA scenarios.
@@ 3.1 Preliminaries
In this paper, we focus on graph test-time adaptation (GTTA) for node classification. A labeled source graph is denoted as ${\mathcal{S}}=({\bm{X}}_{S},{\bm{A}}_{S})$ with node attribute matrix ${\bm{X}}_{S}\in\mathbb{R}^{N\times D}$ and adjacency matrix ${\bm{A}}_{S}\in\{0,1\}^{N\times N}$. The corresponding node label matrix is denoted as ${\bm{Y}}_{S}\in\{0,1\}^{N\times C}$. For a node $v_{i}$, we denote its neighbors as ${\mathbb{N}}(v_{i})$ and node degree as $d_{i}$. A GNN model $g_{S}\circ f_{S}(\cdot)$ is pre-trained on the source graph, where $f_{S}$ is the featurizer extracting node-level representations, and $g_{S}$ is the classifier, which is usually a linear layer. The goal of GTTA is to adapt the pre-trained GNN model to enhance node classification accuracy on an unlabeled target graph ${\mathcal{T}}=({\bm{X}}_{T},{\bm{A}}_{T})$ with a different distribution, while the source graph ${\mathcal{S}}$ are not accessible during adaptation. [^1] This setting is also referred to as source-free unsupervised graph domain adaptation (Mao et al., 2024). Here, we primarily follow the terminology used by Jin et al. (2023). It is important to note that, unlike the online setting often adopted in image TTA, graph TTA allows simultaneous access to the entire unlabeled target graph ${\mathcal{T}}$ (Liang et al., 2023).
Compared with TTA on regular data like images, GTTA is more challenging due to the co-existence of attribute shifts and structure shifts (Li et al., 2022a; Wu et al., 2023b), which are formally defined as follows (Liu et al., 2023).
@@
Homophily refers to the phenomenon that a node tends to connect with nodes with the same labels. Formally, the node homophily of a graph $\mathcal{G}$ is defined as (Pei et al., 2020):
$h(\mathcal{G})=\frac{1}{N}\!\sum_{i}h_{i},\quad\text{where }h_{i}=\frac{|\{v_{j}\!\in\!{\mathbb{N}}(v_{i}):y_{j}\!=\!y_{i}\}|}{d_{i}},$ (1)
@@ 3.2 Impacts of distribution shifts
CSBM (Deshpande et al., 2018) is a random graph generator widely used in the analysis of GNNs (Ma et al., 2022; Mao et al., 2023; Yan et al., 2022). Specifically, we consider a CSBM with two classes +=${\mathbb{C}}_{+}=\{v_{i}:y_{i}=+1\}$ and =${\mathbb{C}}_{-}=\{v_{i}:y_{i}=-1\}$, each having $\frac{N}{2}$ nodes. The attributes for each node $v_{i}$ are independently sampled from a Gaussian distribution ${\bm{x}}_{i}\sim{\mathcal{N}}({\bm{\mu}}_{i},{\bm{I}})$, where ${\bm{\mu}}_{i}={\bm{\mu}}_{+}$ for $v_{i}\in{\mathbb{C}}_{+}$ and ${\bm{\mu}}_{i}={\bm{\mu}}_{-}$ for $v_{i}\in{\mathbb{C}}_{-}$. Each pair of nodes are connected with probability $p$ if they are from the same class, otherwise $q$. As a result, the average degree is $d=\frac{N(p+q)}{2}$ and node homophily is $h=\frac{p}{p+q}$. We denote the graph as $\text{CSBM}({\bm{\mu}}_{+},{\bm{\mu}}_{-},d,h)$, where ${\bm{\mu}}_{+},{\bm{\mu}}_{-}$ encode the node attributes and $d,h$ encode the graph structure.
@@
We consider a single-layer GCN, whose featurizer is denoted as ${\bm{Z}}\!=\!f({\bm{X}},\!{\bm{A}})\!=\!{\bm{X}}+\gamma\cdot{\bm{D}}^{-1}{\bm{A}}{\bm{X}}=({\bm{I}}+\gamma\cdot{\bm{D}}^{-1}{\bm{A}}){\bm{X}}$, where ${\bm{D}}$ is the degree matrix. Equivalently, for each node $v_{i}$, its node representation is ${\bm{z}}_{i}={\bm{x}}_{i}+\gamma\cdot\frac{1}{d_{i}}\sum_{v_{j}\in{\mathbb{N}}(v_{i})}{\bm{x}}_{j}$. The parameter $\gamma$ controls the mixture between the node’s own representation and its one-hop neighbors’ average representation. We consider $\gamma$ as a fixed parameter for now, and adapt it later in Subsection 3.3. We consider a linear classifier as $g({\bm{Z}})={\bm{Z}}{\bm{w}}+{\bm{1}}b$, which predicts a node $v_{i}$ as positive if ${\bm{z}}_{i}^{\top}{\bm{w}}+b\geq 0$ and vise versa.
@@
{restatable}
propositionnoderepresentation For graphs generated by $\text{CSBM}({\bm{\mu}}_{+},{\bm{\mu}}_{-},d,h)$, the node representation ${\bm{z}}_{i}$ of node $v_{i}\in{\mathbb{C}}_{+}$ generated by a single-layer GCN follows a Gaussian distribution of
@@
{restatable}
corollarycrlacc When ${\bm{\mu}}_{+}={\bm{\mu}},{\bm{\mu}}_{-}=-{\bm{\mu}}$, and all nodes have the same homophily $h=\frac{p}{p+q}$ and degree $d=\frac{N(p+q)}{2}$, the classifier maximizes the expected accuracy when ${\bm{w}}=\operatorname{sign}(1+\gamma(2h-1))\cdot\frac{{\bm{\mu}}}{\|{\bm{\mu}}\|_{2}}$ and $b=0$. It gives a linear decision boundary of $\{{\bm{z}}:{\bm{z}}^{\top}{\bm{w}}=0\}$ and the expected accuracy
@@
{restatable}
[Impacts of attribute shifts]propositionpropattribute When training a single-layer GCN on a source graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d,h)$, while testing it on a target graph of $\text{CSBM}({\bm{\mu}}+\Delta{\bm{\mu}},-{\bm{\mu}}+\Delta{\bm{\mu}},d,h)$ with $\|\Delta{\bm{\mu}}\|_{2}<|\frac{1+\gamma(2h-1)}{1+\gamma}|\cdot\|{\bm{\mu}}\|_{2}$, we have
@@
{restatable}
[Impacts of structure shifts]propositionpropstructure When training a single-layer GCN on a source graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d_{S},h_{S})$, while testing it on a target graph of $\text{CSBM}({\bm{\mu}},-{\bm{\mu}},d_{T},h_{T})$, where $1\leq d_{T}=d_{S}-\Delta d<d_{S}$ and $\frac{1}{2}<h_{T}=h_{S}-\Delta h<h_{S}$, if $\gamma>0$, we have
@@
Propositions 3.2 and 3.2 imply that attribute shifts and structure shifts impact the accuracy of GNN differently. Specifically, attribute shifts impact the decision boundary of the classifier, while structure shifts significantly degrade the node representation quality. These propositions also match with our empirical findings in Figure 2 and Figure 10 (in Appendix C.1). Since generic TTA methods (Wang et al., 2021; Iwasawa & Matsuo, 2021; Zhang et al., 2023) usually rely on the representation quality and refine the decision boundary, their effectiveness is limited under structure shifts.